OpenStudy (anonymous):

for a right triangle that has legs a and b hypotenuse c, find the length of the side not given here a=1 c= square root 3

OpenStudy (anonymous):

So by Pythagoras rule a^2 + b^2 =c^2 solve 1^2 + b^2=3 so b^2= 3-1 so b=sqrt(2)

OpenStudy (anonymous):

Let me know and I'll elaborate